# Towards Adaptive Humanoid Control via Multi-Behavior Distillation and Reinforced Fine-Tuning

### v1

@@ Abstract
Humanoid robots are promising to learn a diverse set of human-like locomotion behaviors, including standing up, walking, running, and jumping. However, existing methods predominantly require training independent policies for each skill, yielding behavior-specific controllers that exhibit limited generalization and brittle performance when deployed on irregular terrains and in diverse situations. To address this challenge, we propose Adaptive Humanoid Control (AHC) that adopts a two-stage framework to learn an adaptive humanoid locomotion controller across different skills and terrains. Specifically, we first train several primary locomotion policies and perform a multi-behavior distillation process to obtain a basic multi-behavior controller, facilitating adaptive behavior switching based on the environment. Then, we perform reinforced fine-tuning by collecting online feedback in performing adaptive behaviors on more diverse terrains, enhancing terrain adaptability for the controller. We conduct experiments in both simulation and real-world experiments in Unitree G1 robots. The results show that our method exhibits strong adaptability across various situations and terrains. Project website: https://ahc-humanoid.github.io.
@@ Introduction
Humanoid robots, due to their human-like morphology, are expected to possess various fundamental human-like locomotion abilities, such as walking, running, and standing up after a fall. Previous studies adopt methods such as whole-body control (sentis2006whole), model predictive control (li2023multi), and reinforcement learning (RL) (RL-survey; wang2025more) to enhance the locomotion capabilities of humanoid robots. Recently, RL has achieved a remarkable progress for humanoid locomotion (HumanGym1; HumanGym2; xie2025humanoid), supported by large-scale simulation (isaac; zakka2025mujoco) and advanced policy gradient methods (PPO1; PPO2).
@@
However, existing locomotion algorithms predominantly require training an independent policy with a separately designed reward function, such as standing up (standup1; standup2), jumping (jump), squatting (homie), and walking (HumanGym1). Such a training paradigm yields independent controllers that excel in performing specific skills, while they exhibit limited generalization abilities in both behavior diversity and terrain adaptability. It is noted that directly extending the learning paradigm to multi-skill setting is challenging, which primarily arises from the conflicts of policy gradients caused by different reward functions (yu2020gradient; liu2021conflict), hindering the convergence of multi-skill policies.
Although several works focus on quadruped robot multi-skill learning using Mixture-of-Experts (MoE) or policy distillation (wang2025skill; Yang_2020; huang2025moeloco), multi-skill learning for humanoid robots still remains challenging and requires further exploration (Parkour; long2024).
@@
Training an adaptive multi-behavior controller from scratch via online RL is challenging, as diverse behaviors require distinct environment settings and reward functions, often leading to poor policy convergence (sodhani2021multitaskrl; liu2021conflict). This is primarily caused by issues such as gradient conflict and gradient imbalance. To address this issue and enable the policy to learn multiple primary behaviors, we first train two basic behavior-specific policies using Proximal Policy Optimization (PPO) (PPO1), which are then distilled into a basic multi-behavior policy $\pi^{d}$.
@@
Recovering from unexpected falls is an essential capability for ensuring the robustness and autonomy of humanoid robots. HoST (standup2) has shown that a standing-up control policy can be trained via multiple critics and force curriculum. We also employ multiple critics (mysore2022multi) for our basic standing-up policy $\pi^{b}_{r}$. In the surrogate loss for policy gradient, the advantage function is estimated using a weighted formulation $\hat{A}=\sum_{i=0}^{n}\omega_{i}\cdot(\hat{A}_{i}-\mu_{\hat{A}_{i}})/\sigma_{\hat{A}_{i}}$, where $\omega_{i}$ is a weighting coefficient, and $\mu_{\hat{A}_{i}}$ and $\sigma_{\hat{A}_{i}}$ correspond to the batch mean and standard deviation of the advantage from group $i$, respectively.
Different from HoST, we let the policy focus on recovery on flat terrains. Specifically, the robot is initialized in either a supine or prone position, with additional joint initialization noise introduced to encourage learning robust recovery behaviors from various postures. To mitigate the negative impact caused by interference between sampled rollouts from different initial postures (standup2) and to promote more natural standing-up motions, we introduce AMP-based (AMP; escontrela2022adversarial) reward function using a discriminator which determines whether an episode is a positive sample (i.e., from the reference motion) or a negative sample (i.e., from the robot). The output of the discriminator can be used to guide the robot to recover in a smooth and plausible manner. Consequently, the standing-up policy $\pi^{b}_{r}$ can robustly recover from diverse abnormal postures and learns behaviors from the reference motion, such as leveraging the arms to support the ground during the standing-up process. The AMP-based reward formulation and the objective of the discriminator can be found in Appendix A.
@@
Locomotion across diverse terrains is also essential for humanoid robots. Although complex system designs and multi-stage training pipelines can enable robots to learn diverse locomotion skills (hwcloco; Parkour), we only adopt a simple framework and reward functions design to enable the policy $\pi^{b}_{w}$ to learn fundamental locomotion ability.
@@
In the behavior distillation process, we use DAgger (chen2020learning; ross2011reduction) to distill different behavior knowledge into an MoE-based multi-behavior policy $\pi^{d}$, which eliminates the gradient conflicts resulting from different reward landscapes of various behaviors.
@@
To prevent tasks with larger reward scales from dominating gradient updates and hindering others’ learning (chen2018gradnorm; hessel2019multi), we use behavior-specific critics with a shared actor during fine-tuning. By assigning a separate critic to each task, we isolate value function learning for task-specific reward functions which enables more accurate value estimation and allows for customized critic architectures for each task (e.g., multiple critics for standing-up behavior). Meanwhile, a shared actor is updated using policy gradients aggregated across tasks, allowing skill transfer and terrain adaptability.
@@
To alleviate this issue, we apply Projecting Conflicting Gradients (PCGrad) (yu2020gradient) to resolve gradient conflicts during optimization. For each pair of task gradients, if they conflict (i.e., their cosine similarity is negative), the gradient of one task is projected onto the normal plane of the other, removing the conflicting direction while preserving progress on the remaining subspace. Specifically, given two task gradients $\mathbf{g}_{i}$ and $\mathbf{g}_{j}$, the projected gradient is computed as:
@@
Following the previous work (rudin2022learning), we adopt the terrain curriculum to improve learning efficiency and adaptability on diverse terrains. An automatic level adjustment mechanism modulates terrain difficulty based on task-specific performance. We design flat, slope, hurdle and discrete terrains for both tasks. The maximum slope inclination is 16.6°. Hurdle terrains are composed of regularly spaced vertical obstacles with a maximum height 0.1mm. The discrete terrain consists of randomly placed rectangular blocks. Specifically, we generate 20 rectangular obstacles with width and lengths sampled between 0.5mm and 2.0mm, and heights uniformly sampled between 0.03mm and 0.15mm. We arrange the terrain map into a 10 × 12 grid of $\times 8mm$ patches, with 10 difficulty levels and 3 columns per terrain type.
@@
HOMIE (homie) 0.802 6.421 0.599 4.795 0.407 3.259 0.442 3.603 - - - -
HoST (standup2) - - - - 0.997 0.978 0.911 0.843
@@
To evaluate the terrain-adaptive capability of our proposed AHC framework, we compare it with the following methods: (1) HOMIE (homie): We adapt the lower-body locomotion policy from HOMIE to our setting by re-training it on our terrain settings. (2) HoST (standup2): We train the standing-up controller from HoST on our terrain settings. (3) Fail Recovery Policy $\pi_{r}^{b}$: The fail recovery behavior policy trained in the first stage of our framework. (4) Walking Policy $\pi_{w}^{b}$: The walking behavior policy trained in the first stage of our framework. (5) Basic Multi-Behaviors Policy $\pi^{d}$: The basic multi-behavior policy from the first stage distillation process.
@@
We compare AHC, which incorporates AMP, with the HoST (standup2) baseline without human motion priors. As shown in Figure 3, we visualize a sequence of snapshots during the stand-up process under both approaches in lying and prone posture. Without AMP (i.e., HoST), the robot exhibits uncoordinated and jerky motions, relying on abrupt limb movements to return to a standing posture. In contrast, AHC policy generates a natural get-up motion, including leg folding, arm support, and trunk lifting. We further compare the joint velocity accelerate during the recovery. As shown in Figure 4, the velocity curves of key joints (i.e., hip and knee) for AHC policy exhibit fewer abrupt fluctuations compared to the policy trained with HoST. These results demonstrate that AMP helps shape the recovery controller towards producing stable motions, which are difficult to obtain with handcrafted reward functions.
@@
Deep reinforcement learning (RL) algorithms have enabled humanoid robots to perform robust and even highly complex locomotion behaviors in simulation environments (isaac; todorov2012mujoco). Without relying on external sensors, prior works have demonstrated several challenging behaviors, including coordinated control of the upper and lower body (almi), whole-body locomotion (2024realrlloco), robust traversal over complex terrains (gu2024advancinghumanoidlocomotionmastering), and highly flexible full-body teleoperation (homie). Fundamental capabilities such as fall recovery have also been achieved through techniques like using multiple critics (standup2). The RL policy with external sensors like depth cameras and LiDARs is able to perceive the environment, like terrains and obstacles, and learn to avoid obstacles and leap a gap (pim; vb-com). More extreme terrains with sparse footholds can be traversed by employing finer terrain perception or attention-based network designs (wang2025beamdojo; he2025attentionbasedmapencodinglearning). However, all these works focus solely on a single behavior, such as moving or recovery. In contrast, our proposed framework enables the robot to acquire multiple skills and autonomously select appropriate behaviors based on its current state.
@@
Mastering multiple behaviors is essential for enhancing the adaptability and practical application of robots. Policy distillation allows the integration of skills from multiple expert policies into a single policy, enabling diverse behaviors for navigating complex terrains (robotparkour; Parkour). Alternatively, hierarchical frameworks can select among multiple skill policies to facilitate efficient multi-skill traversal (Anymalparkour). HugWBC (xue2025unified) leverages input signals such as gait frequency and foot contact patterns to guide the policy, allowing it to exhibit different behaviors in response to varying commands. MoE-Loco (huang2025moeloco) adopts an MoE architecture to reduce gradient conflicts in multi-skill RL, thereby improving training efficiency. MoRE (wang2025more) further enhances policy performance by incorporating AMP-based rewards and external sensor inputs. However, prior works typically achieve multi-behavior capabilities either through explicit control signals or by combining behaviors with high similarity (e.g., stair climbing and gap jumping). In contrast, our proposed framework integrates highly diverse behaviors into a single unified policy and enables the robot to autonomously switch between them based on its state.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No.62306242), the Young Elite Scientists Sponsorship Program by CAST (Grant No. 2024QNRC001), and the Yangfan Project of the Shanghai (Grant No.23YF11462200).
@@ A. AMP Reward Formulation and Discriminator Objective
We adopt Adversarial Motion Prior (AMP) (AMP; escontrela2022adversarial) to provide a style reward that encourages natural behaviors. The AMP consists of a discriminator $D_{\phi}$ that determines whether a sequence of states originates from the reference motion dataset or from the policy. We construct the AMP input state $s_{t}^{\text{amp}}\in\mathbb{R}^{20}$ by extracting 20 joint positions from the full observation. Unlike previous works that rely on only two consecutive states, we provide the discriminator with a temporal context by feeding a 5-step window of AMP states. Specifically, the input sequence of discriminator is defined as $\tau_{t}=(s_{t-3}^{\text{amp}},s_{t-2}^{\text{amp}},s_{t-1}^{\text{amp}},s_{t}^{\text{amp}},s_{t+1}^{\text{amp}})$.
@@ B. Training Details
Our PPO implementation follows the standard formulation with clipped surrogate loss and generalized advantage estimation (GAE). We use Adam optimizer with the learning rate of $1\times 10^{-3}$. Each PPO iteration collects rollouts of 32 environment steps, followed by 5 learning epoch using 4 minibatches. We set the discount factory $\gamma=0.99$, GAE lambda $\lambda=0.95$, clipping ratio 0.2, and value loss coefficient 1.0. The reward definitions for both the recovery and locomotion tasks are detailed in Table 3. Following the implementation of (standup2), the rewards for the recovery task are also grouped into four categories.
@@
To alleviate gradient interference between different task objectives during fine-tuning, we employ a gradient surgery strategy following the procedure described in (yu2020gradient). Specifically, after computing the task-specific policy gradients $\bm{g}_{w}$ and $\bm{g}_{r}$ from the walk and recovery objectives, respectively, we project one gradient onto the normal plane of the other whenever a conflicting direction is detected (i.e., when $\langle\bm{g}_{w},\bm{g}_{r}\rangle<0$). Since our framework involves only two task, we randomly select the projection direction at each update step, i.e., either projecting the locomotion gradient onto the recovery gradient or vice versa. This stochastic projection prevents bias toward any single task and maintains balanced optimization across behaviors.
Both training stages use the same domain randomization settings and joint PD gains ($K_{p}$ and $K_{d}$), as shown in Table 4 and Table 5 respectively. The domain randomization terms follow the configurations commonly adopted in previous works (rudin2022learning; standup1), while the ranges of randomized parameters are further adjusted based on empirical performance observed in real-world experiments. The randomization parameters are resampled at the beginning of each episode to prevent overfitting to specific simulation dynamics and to encourage policy robustness across varying environmental and physical conditions.
Table 3: Reward Functions Used in Both Training Stages. The $f_{\text{tol}}$ adopts a Gaussian-style formulation, as detailed in (standup2).

### v2

@@ Abstract
Humanoid robots are promising to learn a diverse set of human-like locomotion behaviors, including standing up, walking, running, and jumping. However, existing methods predominantly require training independent policies for each skill, yielding behavior-specific controllers that exhibit limited generalization and brittle performance when deployed on irregular terrains and in diverse situations. To address this challenge, we propose Adaptive Humanoid Control (AHC) that adopts a two-stage framework to learn an adaptive humanoid locomotion controller across different skills and terrains. Specifically, we first train several primary locomotion policies and perform a multi-behavior distillation process to obtain a basic multi-behavior controller, facilitating adaptive behavior switching based on the environment. Then, we perform reinforced fine-tuning by collecting online feedback in performing adaptive behaviors on more diverse terrains, enhancing terrain adaptability for the controller. We conduct experiments in both simulation and real-world experiments in Unitree G1 robots. The results show that our method exhibits strong adaptability across various situations and terrains.
Websitehttps://ahc-humanoid.github.io
@@ Introduction
Humanoid robots, due to their human-like morphology, are expected to possess various fundamental human-like locomotion abilities, such as walking, running, and standing up after a fall. Previous studies adopt methods such as whole-body control (Sentis and Khatib 2006), model predictive control (Li and Nguyen 2023), and reinforcement learning (RL) (Ernst and Louette 2024; Wang et al. 2025b) to enhance the locomotion capabilities of humanoid robots. Recently, RL has achieved a remarkable progress for humanoid locomotion (Gu, Wang, and Chen 2024; Gu et al. 2024a; Xie et al. 2025), supported by large-scale simulation (Makoviychuk et al. 2021; Zakka et al. 2025) and advanced policy gradient methods (Schulman et al. 2017; Engstrom et al. 2019).
@@
However, existing locomotion algorithms predominantly require training an independent policy with a separately designed reward function, such as standing up (He et al. 2025b; Huang et al. 2025b), jumping (Tan et al. 2024), squatting (Ben et al. 2025), and walking (Gu, Wang, and Chen 2024). Such a training paradigm yields independent controllers that excel in performing specific skills, while they exhibit limited generalization abilities in both behavior diversity and terrain adaptability. It is noted that directly extending the learning paradigm to multi-skill setting is challenging, which primarily arises from the conflicts of policy gradients caused by different reward functions (Yu et al. 2020; Liu et al. 2021), hindering the convergence of multi-skill policies.
Although several works focus on quadruped robot multi-skill learning using Mixture-of-Experts (MoE) or policy distillation (Wang et al. 2025a; Yang et al. 2020; Huang et al. 2025a), multi-skill learning for humanoid robots still remains challenging and requires further exploration (Zhuang, Yao, and Zhao 2024; Long et al. 2024a).
@@
Training an adaptive multi-behavior controller from scratch via online RL is challenging, as diverse behaviors require distinct environment settings and reward functions, often leading to poor policy convergence (Sodhani, Zhang, and Pineau 2021; Liu et al. 2021). This is primarily caused by issues such as gradient conflict and gradient imbalance. To address this issue and enable the policy to learn multiple primary behaviors, we first train two basic behavior-specific policies using Proximal Policy Optimization (PPO) (Schulman et al. 2017), which are then distilled into a basic multi-behavior policy $\pi^{d}$.
@@
Recovering from unexpected falls is an essential capability for ensuring the robustness and autonomy of humanoid robots. HoST (Huang et al. 2025b) has shown that a standing-up control policy can be trained via multiple critics and force curriculum. We also employ multiple critics (Mysore et al. 2022) for our basic standing-up policy $\pi^{b}_{r}$. In the surrogate loss for policy gradient, the advantage function is estimated using a weighted formulation $\hat{A}=\sum_{i=0}^{n}\omega_{i}\cdot(\hat{A}_{i}-\mu_{\hat{A}_{i}})/\sigma_{\hat{A}_{i}}$, where $\omega_{i}$ is a weighting coefficient, and $\mu_{\hat{A}_{i}}$ and $\sigma_{\hat{A}_{i}}$ correspond to the batch mean and standard deviation of the advantage from group $i$, respectively.
Different from HoST, we let the policy focus on recovery on flat terrains. Specifically, the robot is initialized in either a supine or prone position, with additional joint initialization noise introduced to encourage learning robust recovery behaviors from various postures. To mitigate the negative impact caused by interference between sampled rollouts from different initial postures (Huang et al. 2025b) and to promote more natural standing-up motions, we introduce AMP-based (Peng et al. 2021; Escontrela et al. 2022) reward function using a discriminator which determines whether an episode is a positive sample (i.e., from the reference motion) or a negative sample (i.e., from the robot). The output of the discriminator can be used to guide the robot to recover in a smooth and plausible manner. Consequently, the standing-up policy $\pi^{b}_{r}$ can robustly recover from diverse abnormal postures and learns behaviors from the reference motion, such as leveraging the arms to support the ground during the standing-up process. The AMP-based reward formulation and the objective of the discriminator can be found in Appendix A.
@@
Locomotion across diverse terrains is also essential for humanoid robots. Although complex system designs and multi-stage training pipelines can enable robots to learn diverse locomotion skills (Lin et al. 2025; Zhuang, Yao, and Zhao 2024), we only adopt a simple framework and reward functions design to enable the policy $\pi^{b}_{w}$ to learn fundamental locomotion ability.
@@
In the behavior distillation process, we use DAgger (Chen et al. 2020; Ross, Gordon, and Bagnell 2011) to distill different behavior knowledge into an MoE-based multi-behavior policy $\pi^{d}$, which eliminates the gradient conflicts resulting from different reward landscapes of various behaviors.
@@
To prevent tasks with larger reward scales from dominating gradient updates and hindering others’ learning (Chen et al. 2018; Hessel et al. 2019), we use behavior-specific critics with a shared actor during fine-tuning. By assigning a separate critic to each task, we isolate value function learning for task-specific reward functions which enables more accurate value estimation and allows for customized critic architectures for each task (e.g., multiple critics for standing-up behavior). Meanwhile, a shared actor is updated using policy gradients aggregated across tasks, allowing skill transfer and terrain adaptability.
@@
To alleviate this issue, we apply Projecting Conflicting Gradients (PCGrad) (Yu et al. 2020) to resolve gradient conflicts during optimization. For each pair of task gradients, if they conflict (i.e., their cosine similarity is negative), the gradient of one task is projected onto the normal plane of the other, removing the conflicting direction while preserving progress on the remaining subspace. Specifically, given two task gradients $\mathbf{g}_{i}$ and $\mathbf{g}_{j}$, the projected gradient is computed as:
@@
Following the previous work (Rudin et al. 2022), we adopt the terrain curriculum to improve learning efficiency and adaptability on diverse terrains. An automatic level adjustment mechanism modulates terrain difficulty based on task-specific performance. We design flat, slope, hurdle and discrete terrains for both tasks. The maximum slope inclination is 16.6°. Hurdle terrains are composed of regularly spaced vertical obstacles with a maximum height 0.1mm. The discrete terrain consists of randomly placed rectangular blocks. Specifically, we generate 20 rectangular obstacles with width and lengths sampled between 0.5mm and 2.0mm, and heights uniformly sampled between 0.03mm and 0.15mm. We arrange the terrain map into a 10 × 12 grid of $\times 8mm$ patches, with 10 difficulty levels and 3 columns per terrain type.
@@
HOMIE (Ben et al. 2025) 0.802 6.421 0.599 4.795 0.407 3.259 0.442 3.603 - - - -
HoST (Huang et al. 2025b) - - - - 0.997 0.978 0.911 0.843
@@
To evaluate the terrain-adaptive capability of our proposed AHC framework, we compare it with the following methods: (1) HOMIE (Ben et al. 2025): We adapt the lower-body locomotion policy from HOMIE to our setting by re-training it on our terrain settings. (2) HoST (Huang et al. 2025b): We train the standing-up controller from HoST on our terrain settings. (3) Fail Recovery Policy $\pi_{r}^{b}$: The fail recovery behavior policy trained in the first stage of our framework. (4) Walking Policy $\pi_{w}^{b}$: The walking behavior policy trained in the first stage of our framework. (5) Basic Multi-Behaviors Policy $\pi^{d}$: The basic multi-behavior policy from the first stage distillation process.
@@
We compare AHC, which incorporates AMP, with the HoST (Huang et al. 2025b) baseline without human motion priors. As shown in Figure 3, we visualize a sequence of snapshots during the stand-up process under both approaches in lying and prone posture. Without AMP (i.e., HoST), the robot exhibits uncoordinated and jerky motions, relying on abrupt limb movements to return to a standing posture. In contrast, AHC policy generates a natural get-up motion, including leg folding, arm support, and trunk lifting. We further compare the joint velocity accelerate during the recovery. As shown in Figure 4, the velocity curves of key joints (i.e., hip and knee) for AHC policy exhibit fewer abrupt fluctuations compared to the policy trained with HoST. These results demonstrate that AMP helps shape the recovery controller towards producing stable motions, which are difficult to obtain with handcrafted reward functions.
@@
Deep reinforcement learning (RL) algorithms have enabled humanoid robots to perform robust and even highly complex locomotion behaviors in simulation environments (Makoviychuk et al. 2021; Todorov, Erez, and Tassa 2012). Without relying on external sensors, prior works have demonstrated several challenging behaviors, including coordinated control of the upper and lower body (Shi et al. 2025), whole-body locomotion (Radosavovic et al. 2024), robust traversal over complex terrains (Gu et al. 2024b), and highly flexible full-body teleoperation (Ben et al. 2025). Fundamental capabilities such as fall recovery have also been achieved through techniques like using multiple critics (Huang et al. 2025b). The RL policy with external sensors like depth cameras and LiDARs is able to perceive the environment, like terrains and obstacles, and learn to avoid obstacles and leap a gap (Long et al. 2024b; Ren et al. 2025). More extreme terrains with sparse footholds can be traversed by employing finer terrain perception or attention-based network designs (Wang et al. 2025c; He et al. 2025a). However, all these works focus solely on a single behavior, such as moving or recovery. In contrast, our proposed framework enables the robot to acquire multiple skills and autonomously select appropriate behaviors based on its current state.
@@
Mastering multiple behaviors is essential for enhancing the adaptability and practical application of robots. Policy distillation allows the integration of skills from multiple expert policies into a single policy, enabling diverse behaviors for navigating complex terrains (Zhuang et al. 2023; Zhuang, Yao, and Zhao 2024). Alternatively, hierarchical frameworks can select among multiple skill policies to facilitate efficient multi-skill traversal (Hoeller et al. 2024). HugWBC (Xue et al. 2025) leverages input signals such as gait frequency and foot contact patterns to guide the policy, allowing it to exhibit different behaviors in response to varying commands. MoE-Loco (Huang et al. 2025a) adopts an MoE architecture to reduce gradient conflicts in multi-skill RL, thereby improving training efficiency. MoRE (Wang et al. 2025b) further enhances policy performance by incorporating AMP-based rewards and external sensor inputs. However, prior works typically achieve multi-behavior capabilities either through explicit control signals or by combining behaviors with high similarity (e.g., stair climbing and gap jumping). In contrast, our proposed framework integrates highly diverse behaviors into a single unified policy and enables the robot to autonomously switch between them based on its state.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No.62306242, No.62302120), the Young Elite Scientists Sponsorship Program by CAST (Grant No.2024QNRC001), the Yangfan Project of the Shanghai (Grant No.23YF11462200), and the Heilongjiang Key R&D Program of China (Grant No.GA23A915).
@@ A. AMP Reward Formulation and Discriminator Objective
We adopt Adversarial Motion Prior (AMP) (Peng et al. 2021; Escontrela et al. 2022) to provide a style reward that encourages natural behaviors. The AMP consists of a discriminator $D_{\phi}$ that determines whether a sequence of states originates from the reference motion dataset or from the policy. We construct the AMP input state $s_{t}^{\text{amp}}\in\mathbb{R}^{20}$ by extracting 20 joint positions from the full observation. Unlike previous works that rely on only two consecutive states, we provide the discriminator with a temporal context by feeding a 5-step window of AMP states. Specifically, the input sequence of discriminator is defined as $\tau_{t}=(s_{t-3}^{\text{amp}},s_{t-2}^{\text{amp}},s_{t-1}^{\text{amp}},s_{t}^{\text{amp}},s_{t+1}^{\text{amp}})$.
@@ B. Training Details
Our PPO implementation follows the standard formulation with clipped surrogate loss and generalized advantage estimation (GAE). We use Adam optimizer with the learning rate of $1\times 10^{-3}$. Each PPO iteration collects rollouts of 32 environment steps, followed by 5 learning epoch using 4 minibatches. We set the discount factory $\gamma=0.99$, GAE lambda $\lambda=0.95$, clipping ratio 0.2, and value loss coefficient 1.0. The reward definitions for both the recovery and locomotion tasks are detailed in Table 3. Following the implementation of (Huang et al. 2025b), the rewards for the recovery task are also grouped into four categories.
@@
To alleviate gradient interference between different task objectives during fine-tuning, we employ a gradient surgery strategy following the procedure described in (Yu et al. 2020). Specifically, after computing the task-specific policy gradients $\bm{g}_{w}$ and $\bm{g}_{r}$ from the walk and recovery objectives, respectively, we project one gradient onto the normal plane of the other whenever a conflicting direction is detected (i.e., when $\langle\bm{g}_{w},\bm{g}_{r}\rangle<0$). Since our framework involves only two task, we randomly select the projection direction at each update step, i.e., either projecting the locomotion gradient onto the recovery gradient or vice versa. This stochastic projection prevents bias toward any single task and maintains balanced optimization across behaviors.
Both training stages use the same domain randomization settings and joint PD gains ($K_{p}$ and $K_{d}$), as shown in Table 4 and Table 5 respectively. The domain randomization terms follow the configurations commonly adopted in previous works (Rudin et al. 2022; He et al. 2025b), while the ranges of randomized parameters are further adjusted based on empirical performance observed in real-world experiments. The randomization parameters are resampled at the beginning of each episode to prevent overfitting to specific simulation dynamics and to encourage policy robustness across varying environmental and physical conditions.
Table 3: Reward Functions Used in Both Training Stages. The $f_{\text{tol}}$ adopts a Gaussian-style formulation, as detailed in (Huang et al. 2025b).